OpenStudy (anonymous):

What is the simplified form of the expression 4c2 + 5c 4c2 + 7c 8c2 + 5c 8c2 + 7c

OpenStudy (anonymous):

\[\sqrt{16c4} - \sqrt{c2} + 3 \sqrt{c2} + \sqrt{9c2}\]

ganeshie8 (ganeshie8):

nice latex !

ganeshie8 (ganeshie8):

simplify each term first

ganeshie8 (ganeshie8):

take first term : \(\sqrt{16c^4}\)

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

tell me this, when u square, which number gives you \(16\) ?

OpenStudy (anonymous):

4 squred is 16

ganeshie8 (ganeshie8):

yes ! so when u take square root of 16, u wil get 4 back

ganeshie8 (ganeshie8):

\(\sqrt{16} = \sqrt{4^2} \) = \( 4\)

ganeshie8 (ganeshie8):

the "square" and "square root" cancel out

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

lets go back tofirst term : \(\sqrt{16c^4}\) \(\sqrt{4^2(c^2)^2}\) \(4c^2\)

ganeshie8 (ganeshie8):

lets do the whole expression :-

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

\(\sqrt{16c4} - \sqrt{c2} + 3 \sqrt{c2} + \sqrt{9c2} \) \( 4c^2 - c + 3 c+ 3c \) simplify

OpenStudy (anonymous):

The answer is 4c2 + 5c

OpenStudy (anonymous):

Whic is the first answer.

ganeshie8 (ganeshie8):

perfect !

OpenStudy (anonymous):

Thank you, you were awesome help.

ganeshie8 (ganeshie8):

np :)